OpenStudy (anonymous):

a number is multiplied by 5. If 5 is added to the product, you get 10. what is the number?

OpenStudy (mathstudent55):

Let x be the number. How do you show the number multiplied by 5?

OpenStudy (anonymous):

idk that's what the problem says

OpenStudy (mathstudent55):

Using x, how do you express 5 times x? It's 5 * x or even simpler, 5x. Ok?

OpenStudy (anonymous):

so the answer is 5x?

OpenStudy (mathstudent55):

No. We are making an equation so we can find the answer. This is just one step in making and equation. A number multiplied by 5 is 5x. Now we add 5 to the product. We add 5 to 5x, so we get 5x + 5

OpenStudy (anonymous):

and?

OpenStudy (mathstudent55):

Then we are told that after doing the operations, the answer is 10. That means 5x + 5 = 10 This is the equation we must solve for x.

OpenStudy (mathstudent55):

5x + 5 = 10 We want to isolate x. Since 5 is being added to 5x, we first subtract 5 from both sides. 5x + 5 = 10 -5 -5 ------------- 5x = 5 Now x is being multiplied by 5. We want x alone, so we divide both sides by 5: 5x/5 = 5/5 x = 1 The answer is: the number is 1

OpenStudy (mathstudent55):

Now we can check to see if the answer is correct. We go back to the original problem: "A number is multiplied by 5. If 5 is added to the product, you get 10. what is the number?" We calculated the number to be 1, so let;'s see if it really works. "A number is multiplied by 5. 5 * 1 = 5 If 5 is added to the product, you get 10. 5 + 5 = 10 Yes, starting with 1, we do get 10, so 1 is the correct answer.